A youth breakfast held a pancake breakfast to raise money. Tickets were $2 for children & $6.50 for adults. If 268 tickets were sold and the group took in $1508, find the number for adults and the number of children that attended the breakfast.
Answer by addingup(3677)

children: c = 2
adults: a = 6.50
c+a = 268 thus c = 268-a
2c+6.50a = 1508
substitute for c
2(268-a)+6.50a = 1508
536-2a+6.50a = 1508
4.50a = 972
a = 216 adult tickets were sold
268-216 = 52 children